# CORSIKA 8 with Pythia 8: Simulating Vertical Proton Showers

### v1

@@ 1 Pythia 8
In Fig.1, the cross-sections for proton-proton and proton-air interactions, from $\sim$100 GeV to 100 EeV are displayed [^3] Further collisions systems, such as $\pi$p and $\pi$-air, can be found in Gaudu, (2024).. Pythia 8 is overall compatible with the commonly high energy models used in air shower physics: EPOS-LHC, Sibyll 2.3d and QGSJet-II.04. In the proton-proton case, all hadronic interaction models yield an extremely similar description, except for the highest energies where Pythia diverges, while in proton-air interactions, this divergence starts at the TeV scale already, establishing Pythia 8 as the model with the highest cross-sections.
@@ 2 Corsika 8 + Pythia 8
The implementation of Pythia in Corsika 8 follows in the footsteps of previous studies by Reininghaus, (2024) and Reininghaus et al., (2023), but now using the Angantyr model in Pythia 8.312 and its opportunities as discussed in Sjöstrand and Utheim, (2022) and Gaudu, (2024). A new feature allows for a single instance of Pythia to run, from which the beam system can be switched and the collision energy can vary on an event-by-event basis.
@@
For this work, 300 vertical air showers that were induced by protons with a primary energy of 1017 eV were simulated using four different high-energy interaction models. The following energy thresholds for particle tracking were used: e±/$\gamma$ particles at 10 MeV, hadrons and muons at 300 MeV. Electromagnetic particles with an energy below 10-6 times the primary energy were combined statistically to reduce runtime; this thinning is detailed in Alameddine et al., (2023). For technical reasons, the energy threshold to transition from the lower energy model FLUKA (The FLUKA Collaboration et al., (2024)) to Pythia 8 is currently set at 4 TeV, while the Corsika 8 default value is $\sim$80 GeV when using the other high energy models.
@@ 2.1 Shower development
The longitudinal profile of Pythia 8 showers for several particle types (photons, electrons/positrons, hadrons and muons) are displayed in Fig. 2. These profiles are compatible to those from Alameddine et al., (2023) and (cite C8@ECRS2024) when previously comparing Corsika 8 to Corsika 7 outputs, with interaction models EPOS-LHC, Sibyll 2.3d and QGSJet-II.04. All models display a crossing behaviour relative to EPOS-LHC, with higher muon content predicted in the early stages of shower development, and similar or lower content at later stages. Pythia 8 and QGSJet-II.04 predict up to 50% more muons at the early stages but end up with about 10% fewer muons by $X\sim 1000$ g cm-2 compared to EPOS-LHC.
@@ Conclusion
The strengh of Pythia resides in its tuning opportunities, allowing in a friendly and well-documented manner its users to modify internal parameters of the model with a strong impact on air shower observables. Therefore, once the investigations of vertical showers with nuclear primaries and inclined air showers for both proton and nuclear primaries are finalized, a tuning study of Pythia will be the next logical step towards refining Pythia in relation to the Muon Puzzle. A forward physics tune of Pythia was previously presented by Fieg et al., (2024), while ongoing efforts for a tune using fixed-target measurement are discussed by Gaudu, (2024).
@@ Acknowledgements
C. Gaudu acknowledges funding from the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) via the Collaborative Research Center SFB1491: Cosmic Interacting Matters – from Source to Signal (F4) – project no. 445052434.
F. Riehn received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101065027.

### v2

@@ 1 Pythia 8
In Fig.1, the cross-sections for proton-proton and proton-air interactions, from $\sim$100 GeV to 100 EeV are displayed [^3] Further collisions systems, such as $\pi$p and $\pi$-air, can be found in Gaudu, 2024a .. Pythia 8 is overall compatible with the commonly high energy models used in air shower physics: EPOS-LHC, Sibyll 2.3d and QGSJet-II.04. In the proton-proton case, all hadronic interaction models yield an extremely similar description, except for the highest energies where Pythia diverges, while in proton-air interactions, this divergence starts at the TeV scale already, establishing Pythia 8 as the model with the highest cross-sections.
@@ 2 Corsika 8 + Pythia 8
The implementation of Pythia in Corsika 8 follows in the footsteps of previous studies by Reininghaus, (2024) and Reininghaus et al., (2023), but now using the Angantyr model in Pythia 8.312 and its opportunities as discussed in Sjöstrand and Utheim, (2022) and Gaudu, 2024a . A new feature allows for a single instance of Pythia to run, from which the beam system can be switched and the collision energy can vary on an event-by-event basis.
@@
For this work, 300 vertical air showers that were induced by protons with a primary energy of 1017 eV were simulated using four different high-energy interaction models. The following energy thresholds for particle tracking were used: e±/$\gamma$ particles at 10 MeV, hadrons and muons at 300 MeV. Electromagnetic particles with an energy below 10-6 times the primary energy were combined statistically to reduce runtime; this thinning is detailed in Alameddine et al., (2023). For technical reasons, the energy threshold to transition from the lower energy model FLUKA (Ballarini et al., (2024)) to Pythia 8 is currently set at 4 TeV, while the Corsika 8 default value is $\sim$80 GeV when using the other high energy models.
@@ 2.1 Shower development
The longitudinal profile of Pythia 8 showers for several particle types (photons, electrons/positrons, hadrons and muons) are displayed in Fig. 2. These profiles are compatible to those from Alameddine et al., (2023) and Gaudu, 2024b when previously comparing Corsika 8 to Corsika 7 outputs, with interaction models EPOS-LHC, Sibyll 2.3d and QGSJet-II.04. All models display a crossing behaviour relative to EPOS-LHC, with higher muon content predicted in the early stages of shower development, and similar or lower content at later stages. Pythia 8 and QGSJet-II.04 predict up to 50% more muons at the early stages but end up with about 10% fewer muons by $X\sim 1000$ g cm-2 compared to EPOS-LHC.
@@ Conclusion
The strengh of Pythia resides in its tuning opportunities, allowing in a friendly and well-documented manner its users to modify internal parameters of the model with a strong impact on air shower observables. Therefore, once the investigations of vertical showers with nuclear primaries and inclined air showers for both proton and nuclear primaries are finalized, a tuning study of Pythia will be the next logical step towards refining Pythia in relation to the Muon Puzzle. A forward physics tune of Pythia was previously presented by Fieg et al., (2024), while ongoing efforts for a tune using fixed-target measurement are discussed by Gaudu, 2024a .
@@ Acknowledgements
C. Gaudu acknowledges funding from the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) via the Collaborative Research Center SFB1491: Cosmic Interacting Matters – from Source to Signal (F4) – project no. 445052434. F. Riehn received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101065027. We thank T. Sjöstrand and the Pythia 8 collaborators for their support in the implementation of Pythia 8 in Corsika 8.